Benzene and chloroform are toxic solvents that are soluble in each other. In a solution of benzene and chloroform, the mole fraction of benzene is 0.450. What is the mass percent of benzene in the mixture?
1
Expert's answer
2019-10-14T14:30:09-0400

n(C6H6)=0.45 mol

n(CHCl3)=0.55 mol

m(CHCl3)=119.4(g/mol)*0.55mol=65.67g

m(C6H6)=0.45 mol*78(g/mol)=35.1g

w(C6H6)=35.1/35.1+65.67=0.3483=34.83%